OpenStudy (anonymous):

differentiate f(x)=(x+x^(-1))^3

OpenStudy (anonymous):

always expand the brackets first and remember to bring the x^-1 down.

OpenStudy (anonymous):

do that first, and tell me what u get.

OpenStudy (anonymous):

x^-1=1/x right

OpenStudy (anonymous):

\[f(x) = (x + x ^{-1})^{3}\]

OpenStudy (anonymous):

that is your equation and u need to bring the x^-1 down as a denominator to make the power positive and sorry, yes now I understand ur working, x^-1=1/x is correct.